Testing LBC with BP10

Plan is to run some science observations to test LBC guiding under BP10. Two fields: LQG, SN09. OBs in home/LBTO/Calib_OBs/ASTRONtests

  • OBs to run for these tests:
    • LQGr/lqg2.ob and lqg1Y.ob
      • lqg2 looks like it has bright stars on the AO and guide chips. lqg1Y does not have a bright star on the AO chip but does on the guide chip.
    • SN09/

01:50 SY sky flats. Notice Blue filter wheel is not centered --- new LBC build used old configuration file which didn't take into account the changes needed in January 2011 to the encoder positions for Wheel 1. Edited positions and restarted flats.

02:00 SY - Y has lots of stars. S ok. Need about 20 sec to get 10K counts. Scaling is too much - S has twice as many counts as Y.

02:14 VI flats --- this field has a lot of stars. Changed to 12h 30d. Slew was still going on before we got to target. See 0220225/021928 B/R. I loaded one OB to do the slew. And then I stopped this OB mid-slew and started a new one.

02:24 VI flats with 15 sec scaling. 8K counts and stars (but fewer than before). Pause, change scaling to 30sec and restart. OK. 022712/3

02:29 BR flats. B 4000 cnts, R 3000 cnts.

02:45 Slew to ACT0239 for pointing and initial FPIA.

02:50 FPIA - need to give 0.5mm Z global offset. (global offset first was not correct, pupil about same size, but later 1.3mm).

03:00 Noticed PSF left gui turned red. VXWorks rebooting by itself. SX panicked.

03:17 Now running FPIA with DX Global Offset updated to 1.3mm. 031912 B... now FPIA is fitting the pupils. Stopped because a preset was sent by Norm in the midst of this.

03:42 Slewing back to ACT0239. In the meantime, SW group tested some things, including wrap.

03:45 dofpia - converged in 6 iterations (but last Red had large astig suddenly appeared).

04:00 ACT0239 observations 040115/040113 for lbcrangebal. binocular = yes. Repeat. 040552/040549. Blue 040552 does not come to /newdata --- it has been 30 seconds at least... play button is black, so I am redoing the OB. 041052/041040. One more because the last pair do not look so well copointed. 041438/041434. Ran lbcrangebal and offsets were still -4, -3 (CE, CA). 41725/41720.

04:18 Slew to lqg2 - checked pointings and we should have guide and AO stars (barely AO stars...). DOFPIA here. Only 2-3 pupils and the seeing estimates are large and discrepant. One blue pupil was not fit well and returned a large, -1000 nm astig.

04:39 Michelle stopped/restarted PCS. Slewed to field with more pupils 5' south to focus/collimate. dofpia here finds more pupils.

04:44 lqg2.ob is being run (from LQGr/ directory). Estimated seeing was 0.9 R/1.2 B. 044631/044618. Run iqplot on these...

05:00 RPC error on Red trackers. Try to replay the OB... (occurred on 045727)... Looks like the Blue is restarting (050307) but the Red is not...

05:05 Stopped LBC. But Trackers hung at 'uninitialization'. Ran lbckill/lbcstart and it restarted... Red trackers still not coming up. Edited OB to do only the last 2 positions: lqg2end.ob...

05:20 2nd power off/on worked. Running lqg2focus and DOFPIA here. 052310/052315 ... Both blue and red sent large -ve z4, but other Zernikes are small. Converged on Red/Blue in 2 iterations.

05:25 052736 B 6.0 (1.3") /052727 R 5.2 (1.2") There is a focus gradient on Red from bottom to top!

05:36 lqg1Y now... DOFPIA - large +ve z4

05:45 lqg1Y ob running now... 054741 B 5.8 (1.3") /054732 R 5.6 (1.25")

06:00 We lost focus on Red... FWHM is shooting up... The MV on the Red side was not working until recently. 2nd dither position already lost focus. dofpia --- -5400.

06:27 DOFPIA on SN09nr. Done. Estimated seeing is 1.4"/1.5" on B/R.

06:37 Running ND_oldsn_sn09rn_VIe.ob now... Though I edited the OB to move target coordinates 9.1" E and 21.9" S, the bright star is still on the edge of the tech chip 1 and the guide "star" selected is a galaxy. Blue took one images (063907) but Red is hung (at 063859)...
06:51:09.66 timeout on CCD Controller [src/camera/camera.c:1534]
Trying to replay the OB without powering off/on LBC, but so far it appears to be hung. lbckill.sh and lbcstart.sh. Came up with message about RPC error. Note that there are two files to be uploaded still
tftp -b32768 -o 192.168.0.201 put lbcr.20110324.020745.fits d:\lbcr...

06:56 Turn off/on after RPC error message. Blue Filters error upon turning off.

07:00 Handing over to Doug & Konstantina. Guiding with LBC and BP10 was OK, but we need still to get more mileage on BP10 & LBC. Run on technical night and try SN09 at that time.

Summary

  • 20110324.fwhm output: The airmass range was not great, but in this, the FWHM and Ellip values were good for Blue. For Red the FWHM was good for the first exposure after focus/collim, but the mirror was cooling rapidly and without a sufficiently bright star on the focus chip - the field was sparse and whatever stars were on the chip were not bright enough - red did not keep up.
###   Seeing data for LBC, UT=20110324
###   Run time: Thu Mar 24 07:35:50 UTC 2011
###   UT(hr)  FWswl   IMAGE CAM  FILT   Texp   AIRM   FWHM  Ellip  Np
       4.775   1.33  044631  B    Us  180.28  1.198   7.17  0.040  88
       4.837   1.43  045013  B    Us  180.28  1.191   7.67  0.040  77
       4.897   1.27  045349  B    Us  180.28  1.185   6.78  0.045  90
       4.958   1.17  045728  B    Us  180.28  1.180   6.25  0.066  100
       5.052   1.16  050307  B    Us  180.28  1.172   6.17  0.080  93
       5.460   1.15  052736  B    Us  180.28  1.147   6.01  0.042  91
       5.520   1.18  053113  B    Us  180.28  1.144   6.18  0.031  90
       5.795   1.12  054741  B    Us  180.28  1.128   5.81  0.043  122
       5.855   1.00  055117  B    Us  180.28  1.127   5.16  0.030  129
       5.915   0.99  055454  B    Us  180.28  1.126   5.12  0.031  118
       5.976   0.99  055832  B    Us  180.28  1.125   5.13  0.033  129
       6.180   1.32  061047  B    Us  180.28  1.126   6.83  0.040  112
       6.240   1.29  061423  B    Us  180.28  1.127   6.71  0.033  114
       6.300   1.39  061801  B    Us  180.28  1.128   7.22  0.027  102
       6.360   1.37  062137  B    Us  180.28  1.129   7.11  0.032  117
       6.652   1.53  063907  B    Vb  320.28  1.184   7.52  0.104  99
       4.772   1.28  044618  R    YF  180.23  1.198   5.56  0.033  91
       4.837   1.32  045014  R    YF  180.23  1.191   5.71  0.035  89
       4.897   1.16  045350  R    YF  180.23  1.185   5.01  0.039  98
       5.458   1.24  052727  R    YF  180.23  1.147   5.24  0.055  111
       5.520   1.33  053111  R    YF  180.23  1.144   5.62  0.052  99
       5.792   1.33  054732  R    YF  180.23  1.128   5.60  0.045  97
       5.854   1.70  055115  R    YF  180.23  1.127   7.13  0.079  73
       5.914   2.32  055451  R    YF  180.23  1.126   9.74  0.072  62
       5.974   3.23  055827  R    YF  180.23  1.125  13.54  0.086  62
       6.177   1.42  061038  R    YF  180.23  1.126   5.95  0.047  92
       6.239   1.64  061421  R    YF  180.23  1.127   6.87  0.068  84
       6.300   1.93  061800  R    YF  180.23  1.128   8.10  0.047  74
       6.359   2.20  062132  R    YF  180.23  1.129   9.25  0.102  69

  • How well does co-pointing work? Compare a pair of B/R images taken immediately after running lbcrangebal tonight and on a previous night with A05.
    • BP10: 20110324 B/R images are 41725/41720
    • vs A05: 20110129 B/R images are 024047/024043. B@1040,2920 and R@1080,2918.
  • lbcrangebal did not bring the star to the Rotator Center for Blue, but did for Red. On 20110129, however, lbcrangebal brought the stars to the rotator center on both Blue and Red.
    • 20110129: B024047 star @ 1040,2920 and R024043 star @ 1080,2918.
    • 20110324: B041725 star @ 1043,2944 (RC at 1040,2921) and R041720 star @ 1082,2914 (RC at 1080,2918)
B 041438 star @ 1050,2963.75
B 041725 star @ 1043, 2944
R 041434 star @ 1079, 2914 
R 041720 star @ 1082, 2914

So it looks like lbcrangebal is bringing star to RC on Red side, but not on Blue. Red uses the IE/CA but Blue relies on mirror tip/tilt. Need to look at this further.

24 March noon Looked at log files from lbcrangebal and these do show that on subsequent exposures in between which lbcrangebal was run, the star on the Red detector remained fixed but on the Blue detector it moved.

Bfile (x,y) Rfile (x,y) 

lbcb.20110324.040115.fits  lbcr.20110324.040113.fits                star at B(1081., 3008)           R(1095,2983)

lbcrangebal:       
RED star is dCAdx = 8.5414708319931 and dCEdx = 13.830043095602 arcsec from the red rotator center
BLUE star is dCAsx = 6.7988265627346 arcsec and dCEsx = 20.442699366029 arcsec from reference point

rangebal_040331.log:RED MIRROR is at = -1.356, -1.17, 2.033, -9.085, 30.845
rangebal_040331.log:BLUE MIRROR is at = 0.825, 0.832, 0.736, 38.857, -61.465


missing!                              lbcr.20110324.040549.fits 

lbcb.20110324.041052.fits  lbcr.20110324.041040.fits               star at B(1060., 3001)             R(1080,2915)
lbcrangebal: 
RED star is dCAdx = 0.44552937600416 and dCEdx = 0.50307810041518 arcsec from the red rotator center
BLUE star is dCAsx = 8.5269293573431 arcsec and dCEsx = 16.385611851099 arcsec from reference point

rangebal_041218.log:RED MIRROR is at = -1.566, -1.451, 1.965, -15.575, 36.509
rangebal_041218.log:BLUE MIRROR is at = 1.037, 0.941, 0.663, 41.075, -67.162


lbcb.20110324.041438.fits  lbcr.20110324.041434.fits               star at B(1050., 2964)              R(1078., 2910)
lbcrangebal:
RED star is dCAdx = 0.048691713934143 and dCEdx = -0.63169384751963 arcsec from the red rotator center
BLUE star is dCAsx = 4.5663090324096 arcsec and dCEsx = 8.7716501195918 arcsec from reference point

rangebal_041540.log:RED MIRROR is at = -1.478, -1.273, 1.936, -11.957, 35.111
rangebal_041540.log:BLUE MIRROR is at = 0.952, 0.771, 0.633, 37.352, -65.822


lbcb.20110324.041725       lbcr.20110324.041720.fits              star at B(1043,2944)                R(1082,2916)

                                                                   RC  at  B(1040,2921)               R(1080,29rangebal_011651.lo12)

24 March evening I found a pair of R/B images from 20110128 on which I had run rangebal. These are: 20110128: 075154 B and 075225 R. Log file in /home/LBTO/iraf/LOGs/rangebal_075623.log. The subsequent pair shows the stars well centered on the rotator centers. I hadn't run lbcrangebal on these (075741 B and 075814 R) but did just now and confirm they are centered.

I ran the lbcrangebal on the first pair and get similar results to what is in rangebal_075623.log (see okuhn@obs3/scratch/20110324/LOGs/rangebal_011651.log.

I believe the lbcrangebal script is doing the calculations correctly, as I would expect since nothing in it was changed, but the fact that the stars were not on the rotator centers last night after lbcrangebal was run suggests to me that something in the interaction with the telescope/PSF is not working as before.

/home/LBTO/iraf/LOGs/rangebal_075623.log
   LBTtools -->>  Observe:lbcrangebal.cl  <<-- Fri 07:56:23 28-Jan-2011

Log file will be written to: LOGs/rangebal_075623.log

Unlearn old task parameters:

   Images /newdata/lbcb.20110128.075154.fits and /newdata/lbcr.20110128.075225.fits b
oth found.

Get current pointing model CE and CA:
 Are these current CE and CA correct? (n or N if no, any other character if yes)
   Current CE = -38.81683
   Current CA = -161.4026
Getting position of star on Red image
z1=7. z2=475.
  Click on the Pointing Star...
 Star at pixel coordinates: 1061., 2919.

RED rotator center is at: 1080., 2912.
RED star is dCAdx = 4.3726257990173 and dCEdx = 1.2051156051466 arcsec from the red r
otator center
Getting position of star on Blue image
z1=862. z2=1074.
  Click on the Pointing Star...
 Star at pixel coordinates: 994., 2936.

BLUE rotator center is at: 1040., 2921.
BLUE star is dCAsx = 10.549731547874 arcsec and dCEsx = 2.4829781045729 arcsec from r
eference point
RED MIRROR is at = -1.348, -1.065, 0.869, -12.015, 29.99
BLUE MIRROR is at = 0.956, 0.902, 0.677, 35.427, -65.043
TEMPERATURES are = -2.671, -1.404, -2.671, -1.738
Iterating on the range balance algorithm for SX

Iterating on the range balance algorithm for DX

   Moving M1...

obs3.lbt.as.ariz!/scratch/20110324/LOGs/rangebal_011651.log
obs3.lbt.as.ariz!/scratch/20110324/
Observe> !less LOGs/rangebal_011651.log

   LBTtools -->>  Observe:lbcrangebal.cl  <<-- Fri 01:16:51 25-Mar-2011

Log file will be written to: LOGs/rangebal_011651.log

Unlearn old task parameters:

   Images /Repository/20110128/lbcb.20110128.075154.fits and /Repository/20110128/lbcr.20110128.075225.fits both found.

/Repository/20110128/lbcb.20110128.075154.fits[LBCCHIP2] -> 075154.fits
/Repository/20110128/lbcr.20110128.075225.fits[LBCCHIP2] -> 075225.fits
Get current pointing model CE and CA:
 Are these current CE and CA correct? (n or N if no, any other character if yes)
   Current CE = 0.
   Current CA = 0.
Getting position of star on Red image
z1=7. z2=475.
  Click on the Pointing Star...
 Star at pixel coordinates: 1062., 2921.

RED rotator center is at: 1080., 2912.
RED star is dCAdx = 4.1870323082206 and dCEdx = 1.6703414171711 arcsec from the red rotator center
Getting position of star on Blue image
z1=862. z2=1074.
  Click on the Pointing Star...
 Star at pixel coordinates: 996., 2936.

BLUE rotator center is at: 1040., 2921.
BLUE star is dCAsx = 10.103313501936 arcsec and dCEsx = 2.5205936367453 arcsec from reference point
RED MIRROR is at = -1.348, -1.065, 0.869, -12.015, 29.99
BLUE MIRROR is at = 0.956, 0.902, 0.677, 35.427, -65.043
TEMPERATURES are = -2.671, -1.404, -2.671, -1.738
Iterating on the range balance algorithm for SX

Iterating on the range balance algorithm for DX



...Done!

-- OlgaKuhn - 24 Mar 2011
Topic revision: r5 - 25 Mar 2011, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback